x-y=-5
x=3y=3 couldyou please help me understand how to solve this problem?

Answer by stanbon(75887)

x - y =-5
x -3y = 3
----------------
Subtract the 2nd equation from the 1st to solve for "y":
4y = -8
y = -2
----------------
Substitute to solve for "x":
x --2 = -5
x + 2 = -5
x = -7
=================
Cheers,
Stan H.